• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 02:31:11 on December 1,2004

    Entry number 137578

    keyword=Left HRS p0 setting

    After run 4997..

    Cycled the Left Q2 and Q3, then set the Left HRS to p0=2.381 GeV.

    After this... moved the Left HRS to 37.69deg (37.5 floor marker, 64.0 vernier).

    FIGURE 1